Consider the circuit shown at left. At 0.2 seconds
Answer #1

(a) We have relation for after time t

i=\frac{V(1-e^{\frac{-Rt}{L}})}{R}

putting all values

i=\frac{24(1-e^{\frac{-6\times 0.2}{2}})}{6}

i=1.80 A ANS

(b)

Voltage across resistor VR=IR

= 1.80\times6

= 10.8 V

Voltage across inductor VL=(24-10.8) =13.2V ANS

(c)

Energy stored E= (LI2)/2

= (2\times1.82) / 2

= 3.24 J ANS